question: a bag of cashews cost $6.05. there are 4 dozen in each bag. what would the cost of 5 dozen be if the rate is the same?
Answer by stanbon(75887)

question: a bag of cashews cost $6.05. there are 4 dozen in each bag. what would the cost of 5 dozen be if the rate is the same?
---
x/5 = 6.05/4
---
x = (5/4)(6.05)
---
x = $7.56
==============
Cheers,
Stan H.
==============